OpenStudy (anonymous):

Hi there, I want someone to check my result in some questions. there are in integration

OpenStudy (anonymous):

\[(1+x ^{2}) \frac{ dy }{ dx } = \sqrt{32} where y(1) = 2\pi\]

OpenStudy (anonymous):

I have the final answer but I want to confirm it is correct or not!!

OpenStudy (anonymous):

I would help u but i dont know anythin about that type of math.. sorry